OpenStudy (anonymous):

imit as x approaches 2 (f(x)-f(2))/(x-2) when f(x)=4x^2-2x+4

OpenStudy (anonymous):

\(f(2)=4\times 2^2-2\times 2+4=16\)

OpenStudy (anonymous):

so you are after \[\lim_{x\to 2}\frac{4x^2-2x+4-16}{x-2}=\frac{4x^2-2x-12}{x-2}\]

OpenStudy (anonymous):

if you replace \(x\) by \(0\) you get \(\frac{0}{0}\) so you know you can factor and cancel factor as \[\frac{(x-2)(\text{something})}{x-2}=\text{something}\] then replace \(x\) by 2

OpenStudy (anonymous):

rather i meant "if you replace \(x\) by 2"

OpenStudy (anonymous):

can be its means that f'(2) if f(x) = 4x^2-2x+4 so, f'(x) = 8x-2 put x=2, to find f'(2)

OpenStudy (anonymous):

so the limit should = 14

OpenStudy (anonymous):

yes..